OpenStudy (anonymous):

Can someone show me how to simplify (-8x^9/y^-18)^2/3. Please explain step by step. Medal.

OpenStudy (astrophysics):

\[\huge \left( \frac{ x^n }{ y^n } \right)^m \implies \frac{ x^{n \times m} }{ y^{n \times m} }\]

OpenStudy (astrophysics):

Try using that and see what you get

OpenStudy (jhannybean):

First you want tomake everything inside the parenthesis a positive power. Remember, \[\large \frac{1}{x^{-1}} \iff x^1\]

OpenStudy (jhannybean):

So what can you do with the \(\large y^{-18}\)?

OpenStudy (anonymous):

so far I've gotten it down to -8x^6/y^-12

OpenStudy (astrophysics):

\[\frac{ 1 }{ x^{-n} } \implies x^n \implies \frac{ 1 }{ y^{-18} } \implies y^{18}\]

OpenStudy (anonymous):

yeah but it's in a whole set of other things

OpenStudy (astrophysics):

If you can show your work, we can help guide you

OpenStudy (astrophysics):

\[\left( -8x^9y^{18} \right)^{2/3}\]

OpenStudy (astrophysics):

This is what you should have by now, it should be much simpler now.

OpenStudy (anonymous):

okay, based on the first thing you showed me I got it from the original to -8x^9*2/3/y^-18*2/3

OpenStudy (jhannybean):

Trust me, simplify it first.

OpenStudy (anonymous):

that's basically the same thing, just split into the two different parts

OpenStudy (jhannybean):

Because dealing with fractional exponents and negative powers is not pretty.

OpenStudy (astrophysics):

Naw, making it positive is much better, I didn't see your exponent was negative that's why I gave you that rule first, but this is much simpler to deal with.

OpenStudy (jhannybean):

\[\large \left(\frac{-8x^9}{\color{red}{y^{-18}}}\right)^{2/3} \implies (-8x^9\color{red}{y^{18}})^{2/3}\]

OpenStudy (anonymous):

okay, I understand. so from now on when negative exponents I need to try and make things positive so it isn't so messy?

OpenStudy (astrophysics):

Yup

OpenStudy (jhannybean):

Now all you have to do is use the distributive property with your exponent, \(\frac{2}{3}\)

OpenStudy (astrophysics):

Makes it easier to deal with now just distribute the 2/3

OpenStudy (anonymous):

okay let me try really quickly

OpenStudy (jhannybean):

So instead of going through all trouble writing out the function, :\[\large 9\cdot \frac{2}{3} =~?\]\[\large 18 \cdot \frac{2}{3} = ~?\]

OpenStudy (anonymous):

9*2/3=6 18*2/3=12

OpenStudy (jhannybean):

Good, now just replace your simplified powers into the function.

OpenStudy (anonymous):

so now I have -8x^6 y^12

OpenStudy (jhannybean):

And we missed one small part. We also has to distribute the 2/3 to the -8. \[(-8)^{2/3} = ((-8)^2)^{1/3} = 64^{1/3}= ~?\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

it's 4

OpenStudy (jhannybean):

mmhmm. So your final answer would be?

OpenStudy (anonymous):

4x^6 y^12

OpenStudy (jhannybean):

\[\large \checkmark\]

OpenStudy (anonymous):

Thank you for your help! I appreciate it!

OpenStudy (jhannybean):

No problem :)